Calculate the freezing point depression of the given electrolytic solution:

5.7 g CaCl2 (calcium chloride) in 18.6 kg H2O (water)
1
Expert's answer
2020-10-23T09:13:35-0400

 DTf = iKf m

                    DTf = Depression in freezing point

                   Kf = Freezing Point Depression Constant

                m = molality

               i = van’t Hoff factor, for CaCl2 i = 3

Moles of CaCl2 = 5.7 g/110.98 g mol−1 = 0.0513 mol

Mass of water in kg = 18.6 kg

Hence molality of CaCl2 = 0.0513 mol/18.6 kg = 2.758×10-3 mol kg-1

 Kf for water is =1.86 K kg mol–1

 DTf = 3×1.86 K kg mol–1×2.758×10-3 mol kg-1

        = 0.01539 K

Depression in freezing point 0.01539 K